OpenStudy (anonymous):

Factor 7x^2 -1= -50x -8

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you cannot factor it unless everything is on one side of the equal sign

Directrix (directrix):

7x^2 - 1 = -50x - 8 +8 + 8 ------------------------------ 7x^2 + 7 = -50x

Directrix (directrix):

7x^2 + + 7 = -50x + 50x +50x ------------------------------- 7x^2 + 50x + 7 = 0

Directrix (directrix):

@yodawg2015 What are two numbers that multiply to 49 AND also add to 50 ?

OpenStudy (hockeychick23):

7x^2 -1+50x= -50x -8+50x 7x^2+50x-1+8= -8+8 7x^2+50x+7=0 (7x+1)(x+7)